# NIR-Assisted Image Denoising: A Selective Fusion Approach and A Real-World Benchmark Dataset

### v1

@@ Abstract
Despite the significant progress in image denoising, it is still challenging to restore fine-scale details while removing noise, especially in extremely low-light environments. Leveraging near-infrared (NIR) images to assist visible RGB image denoising shows the potential to address this issue, becoming a promising technology. Nonetheless, existing works still struggle with taking advantage of NIR information effectively for real-world image denoising, due to the content inconsistency between NIR-RGB images and the scarcity of real-world paired datasets. To alleviate the problem, we propose an efficient Selective Fusion Module (SFM), which can be plug-and-played into the advanced denoising networks to merge the deep NIR-RGB features. Specifically, we sequentially perform the global and local modulation for NIR and RGB features, and then integrate the two modulated features. Furthermore, we present a Real-world NIR-Assisted Image Denoising (Real-NAID) dataset, which covers diverse scenarios as well as various noise levels. Extensive experiments on both synthetic and our real-world datasets demonstrate that the proposed method achieves better results than state-of-the-art ones. The dataset, codes, and pre-trained models will be publicly available at https://github.com/ronjonxu/NAID.
@@ I Introduction
In low-light conditions, it’s common to use short exposure time and high ISO in imaging to prevent motion blur, while this setting inevitably introduces noise due to the limited number of photons captured by camera. With the development of deep learning [1, 2, 3], many image denoising methods [4, 5, 6, 7, 8, 9, 10] have been proposed to remove the noise. Although great progress has been achieved, it is still challenging for these methods to recover fine-scale details faithfully due to the severely ill-posed nature of denoising. An alternative solution is multi-frame denoising [11, 12, 13, 14, 15, 16], in which multiple successive frames are merged to improve performance. However, it is susceptible to the misalignment between frames, and may be less effective in facing dynamic scenes.
@@
Nevertheless, the inconsistencies between NIR and RGB contents limit the positive effect of NIR images in denoising. Firstly, NIR images are captured under additional NIR light and are monochromatic, which leads to brightness and color discrepancies between the two modalities. Secondly, the NIR images may ‘more-see’ or ‘less-see’ the objects than the visible light ones, primarily due to inherent differences in the optical properties within each spectral domain [21]. For example, as shown in Fig. 1 (a), the RGB image clearly contains textual information, while the corresponding NIR image lacks that. In Fig. 1 (b), the NIR image exhibits extra fruit patterns, while these patterns are absent in the RGB image. DVD [19] and SANet [23] have noticed this problem, but their solutions are complex and less effective.
@@ II-B NIR-assited Image Restoration
Compared to single-image restoration, NIR images have the potential to assist in restoring details from degraded images. The earlier work [31] utilizes gradient constraints for NIR-assisted image denoising. Wang et al. [32] further improves the performance with deep learning methods. SSN [18] proposes a multi-task deep network with state synchronization modules. TC-GAN [33] fuses NIR images and RGB ones based on a texture conditional generative adversarial network. DCMAN [34] employs spatial-temporal-spectral priors to introduce NIR videos for low-light RGB video restoration. However, these methods have overlooked the color and structure inconsistency issues between the NIR images and RGB ones. CCDFuse [35] addresses this issue by combining the local modeling ability of convolutional blocks and the non-local modeling ability of transformer ones to extract local and global features of NIR and RGB images respectively. SANet [23] proposes a guided denoising framework by estimating a clean structure map for the noisy RGB image. Wan et al. [20] disentangle the color and structure components from the NIR images and RGB ones. Besides, a few works [36, 37, 19] incorporate different priors into the network design, like sparse coding [36], deep implicit prior [37] and deep inconsistency prior [19]. However, their intricate designs make it difficult to integrate them into existing advanced restoration networks, hindering their extensions and improvements.s
@@ III Real-NAID: Real-World NIR-Assisted Image Denoising Dataset
Existing publicly available NAID datasets generally lack real-world noisy RGB images paired with clean RGB and NIR images, which limits the investigation in real-world NAID. To break such a limitation, we build a real-world NAID dataset, named Real-NAID. Specifically, we employ high ISO and short exposure time to capture the real-world noisy RGB images, as shown in Fig. 2 (a). For capturing the corresponding clean RGB images, we lower the ISO of the camera and appropriately increase the exposure time, as shown in Fig. 2 (b). To obtain paired NIR images, we activate NIR light to ensure a sufficient supply of NIR illumination and then capture them with a dedicated NIR camera, as shown in Fig. 2 (c).
All images are captured with the Huawei X2381-VG surveillance camera, which is equipped with a built-in NIR illuminator specifically designed for capturing NIR images. To ensure image registration among multiple captures, we carefully position the camera and develop a remote control application to capture images of static objects. When capturing clean RGB images, the camera’s ISO is set to 600. When capturing RGB images with low, middle, and high noise levels, we set ISO to 4000, 12000, and 32000, respectively. It is worth noting that we adjust the exposure time of each noise level to keep the brightness of noisy images relatively constant.
@@ IV-B Selective Fusion Module
Global Modulation Module. GMM should handle the global color and brightness difference between the NIR and RGB ones. As shown in Fig. 4 (a), it takes the current NIR features $\mathbf{F_{N}}$ and the RGB ones $\mathbf{F_{R}}$ as inputs to estimate the NIR global modulation weights $\mathbf{W}^{g}_{\mathbf{N}}$ and the RGB ones $\mathbf{W}^{g}_{\mathbf{R}}$. Detailly, it first concatenates $\mathbf{F_{N}}\in\mathbb{R}^{C\times H\times W}$ and $\mathbf{F_{R}}\in\mathbb{R}^{C\times H\times W}$ as inputs, generating initial weights $\mathbf{W_{init}^{g}}\in\mathbb{R}^{2C\times H\times W}$ by a $1\times 1$ convolutional layer, two $1\times 1$ blocks and another $1\times 1$ convolutional layer. It can be written as,
@@
Local Modulation Module. The Local Modulation Module (LMM) should focus on the structure inconsistency between NIR images and RGB ones. We suggest increasing the receptive field to perceive more structure information from a range of neighboring pixels. In detail, LMM takes the globally modulated NIR features $\mathbf{F}^{g}_{\mathbf{N}}$ and the RGB ones $\mathbf{F}^{g}_{\mathbf{R}}$ as inputs to estimate the local NIR weights $\mathbf{W}^{l}_{\mathbf{N}}$ and the RGB ones $\mathbf{W}^{l}_{\mathbf{R}}$, as shown in Fig. 4 (a). Without complex network design, LMM is built upon GMM by replacing the $1\times 1$ convolutional layer in $1\times 1$ block to a large kernel depth-wise convolutional layer (DWConv) [43] for capturing more local information, as shown in Fig. 4 (c). Finally, $\mathbf{W}^{l}_{\mathbf{N}}$ and $\mathbf{W}^{l}_{\mathbf{R}}$ are employed to get the fused NIR and RGB feature $\mathbf{F_{NR}}$ as,
@@
Methods $\sigma=4$ $\sigma=8$FLOPs (G) Time (ms)
@@
Single-Image Denosing Uformer (CVPR’22) 29.58 / 0.8967 / 0.271 27.36 / 0.8632 / 0.352 19.16 1748
Restormer (CVPR’22) 29.67 / 0.9038 / 0.262 27.41 / 0.8741 / 0.343 70.59 2048
NAFNet (ECCV’22) 29.49 / 0.8959 / 0.263 27.29 / 0.8638 / 0.336 8.10 312
Xformer (ICLR’24) 27.32 / 0.8690 / 0.347 27.32 / 0.8690 / 0.347 73.31 5257
NIR-Assisted Image Denosing FGDNet (TMM’22) 23.91 / 0.8371 / 0.439 22.02 / 0.7374 / 0.436 38.67 479
SANet (CVPR’23) 27.68 / 0.8648 / 0.343 25.28 / 0.8304 / 0.413 161.06 2763
CUNet(TPAMI’20) 28.01 / 0.8558 / 0.332 26.07 / 0.8182 / 0.412 14.48 542
MNNet (IF’22) 28.48 / 0.8994 / 0.274 26.33 / 0.8697 / 0.353 23.68 1360
DVN (AAAI’22) 29.69 / 0.9062 / 0.236 27.43 / 0.8799 / 0.292 104.05 761
NIR-Uformer (Ours) 30.10 / 0.9188 / 0.192 28.03 / 0.9008 / 0.238 24.85 2500
NIR-Restormer (Ours) 30.22 / 0.9209 / 0.193 28.11 / 0.8701 / 0.260 89.17 2747
NIR-NAFNet (Ours) 30.08 / 0.9005 / 0.208 27.86 / 0.8664 / 0.273 13.17 462
NIR-Xformer (Ours) 30.29 / 0.9152 / 0.192 28.09 / 0.8947 / 0.230 95.63 7884
@@ V-A Experimental Settings
Implementation Details. We build our NIR-assisted denoising models by incorporating the proposed SFM into a CNN-based advanced denoising network (i.e., NAFNet [24]) and three Transformer-based ones (i.e., Uformer [8], Restormer [7] and Xformer [10]), which are dubbed NIR-NAFNet, NIR-Uformer, NIR-Restormer, and NIR-Xformer, respectively. All models are trained by the Adam [44] optimizer with $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999 for 120k iterations. The batch size is set to 32 and the patch size is set to $128\times 128$. For synthetic image denoising, the cosine annealing strategy [45] is employed to steadily decrease the learning rate from $2\times 10^{-4}$ to $1\times 10^{-6}$. For real-world image denoising, the initial learning rate is set to $3\times 10^{-4}$ and halved every 20k iterations. All experiments are conducted with PyTorch [46] on an Nvidia GeForce RTX A6000 GPU.
@@ V-B Comparison with State-of-the-Art Methods
Experiments are conducted by comparing our NIR-NAFNet, NIR-Uformer, NIR-Restormer, and NIR-Xformer with 9 models, including 4 single image denoising methods (i.e. NAFNet [24], Uformer [8], Restormer [7], and Xformer [10]) and 5 NIR-assisted image denoising methods (i.e. FGDNet [47], SANet [23], CUNet [36], MNNet [48], and DVN [19]). To evaluate the performance quantitatively, we calculate three metrics on the RGB channels, i.e., Peak Signal to Noise Ratio (PSNR), Structural Similarity (SSIM) [49] and Learned Perceptual Image Patch Similarity (LPIPS) [50]. We also evaluate the inference cost of different models. The #FLOPs when processing a $128\times 128$ patch and the inference time when feeding a $1792\times 1008$ image are reported.
@@
Single-Image Denosing Uformer (CVPR’22) 25.56 / 0.7736 / 0.304 24.52 / 0.7418 / 0.347 23.31 / 0.7091 / 0.389
@@
NIR-Assisted Image Denosing FGDNet (TMM’22) 24.25 / 0.7676 / 0.368 22.89 / 0.7367 / 0.430 21.86 / 0.7080 / 0.509
@@ VI-A Comparisons with NAID Baselines
To demonstrate the effectiveness of our SFM, we compare our methods with NAID baselines using varying backbones. Specifically, we simply sum the NIR and RGB features as our baseline, as shown in Fig. 3 (b). We utilize three Transformer-based denoising networks (i.e., Uformer [8], Restormer [7], Xformer [10]) and a CNN-based one (i.e., NAFNet [24]) to conduct experiments, named ‘Uformer-Baseline’, ‘Restormer-Baseline’, ‘Xformer-Baseline’, and ‘NAFNet-Baseline’, respectively. From Table  IV, it is observed that the performance of NAID baselines can be further improved by integrating our proposed SFM, and the improvement is more significant when the noise level is higher. This is because that simple summation in NAID baselines ignores color and structure discrepancies between NIR-RGB images while the proposed SFM alleviates these issues.
@@
Methods Low-Level Noise Middle-Level Noise High-Level Noise
@@
Spatial-Wised Transformer-Based Uformer-Baseline 25.80 / 0.7904 / 0.269 25.03 / 0.7687 / 0.294 24.02 / 0.7488 / 0.320
NIR-Uformer 25.91 / 0.7917 / 0.276 25.14 / 0.7714 / 0.299 24.28 / 0.7534 / 0.321
Channel-Wised Transformer-Based Restormer-Baseline 26.12 / 0.7948 / 0.265 25.38 / 0.7733 / 0.293 24.61 / 0.7587 / 0.316
NIR-Restormer 26.22 / 0.7963 / 0.265 25.51 / 0.7767 / 0.293 24.76 / 0.7626 / 0.315
Mixed Transformer-Based Xformer-Baseline 26.10 / 0.7927 / 0.264 25.23 / 0.7694 / 0.291 24.41 / 0.7553 / 0.314
NIR-Xformer 26.19 / 0.7959 / 0.264 25.38 / 0.7729 / 0.293 24.66 / 0.7595 / 0.319
CNN-Based NAFNet-Baseline 25.92 / 0.7883 / 0.277 25.00 / 0.7628 / 0.304 24.16 / 0.7429 / 0.329
NIR-NAFNet 26.06 / 0.7905 / 0.274 25.26 / 0.7676 / 0.303 24.48 / 0.7503 / 0.321

### v4

@@ Abstract
Despite the significant progress in image denoising, it is still challenging to restore fine-scale details while removing noise, especially in extremely low-light environments. Leveraging near-infrared (NIR) images to assist visible RGB image denoising shows the potential to address this issue, becoming a promising technology. Nonetheless, existing works still struggle with taking advantage of NIR information effectively for real-world image denoising, due to the content inconsistency between NIR-RGB images and the scarcity of real-world paired datasets. To alleviate the problem, we propose an efficient Selective Fusion Module (SFM), which can be plug-and-played into the advanced denoising networks to merge the deep NIR-RGB features. Specifically, we sequentially perform the global and local modulation for NIR and RGB features, and then integrate the two modulated features. Furthermore, we present a Real-world NIR-Assisted Image Denoising (Real-NAID) dataset, which covers diverse scenarios as well as various noise levels. Extensive experiments on both synthetic and our real-world datasets demonstrate that the proposed method achieves better results than state-of-the-art ones. The dataset, codes, and pre-trained models are available at https://github.com/ronjonxu/NAID.
@@ I Introduction
In low-light conditions, it’s common to use short exposure time and high ISO in imaging to prevent motion blur, while this setting inevitably introduces noise due to the limited number of photons captured by camera. With the development of deep learning [1, 2, 3, 4, 5, 6], many image denoising methods [7, 8, 9, 10, 11, 12, 13] have been proposed to remove the noise. Although great progress has been achieved, it is still challenging for these methods to recover fine-scale details faithfully due to the severely ill-posed nature of denoising. An alternative solution is multi-frame denoising [14, 15, 16, 17, 18, 19], in which multiple successive frames are merged to improve performance. However, it is susceptible to the misalignment between frames, and may be less effective in facing dynamic scenes.
@@
Nevertheless, the color, brightness, and structure inconsistencies between NIR and RGB contents limit the positive effect of NIR images in denoising. First, NIR images are monochromatic, which leads to color discrepancy between the two modalities. Second, NIR images are captured under additional NIR light, and objects exhibit varying spectral reflectance in the RGB (400-720 nm) and NIR band (720-1100 nm). These lead to the brightness inconsistency. Thirdly, the NIR images may ‘less-see’ or ‘more-see’ the objects than the visible light ones, leading to structure inconsistency between the two modalities [24]. On the one hand, different from visible RGB images, NIR images tend to provide information pertinent to material properties rather than the color of objects. An example can be seen in the red box in Fig. 1 (a). As the yellow texts and the background have similar NIR spectral reflectance, the background and texts are indistinguishable in the NIR image while the RGB image clearly contains textual information. On the other hand, NIR has stronger penetrating power than visible light, leading NIR images to ‘more-see’ the objects. As shown in the yellow box in Fig. 1 (b), the NIR image exhibits extra fruit patterns while these patterns are absent in the RGB image. DVD [22] and SANet [26] have noticed this problem, but their solutions are complex and less effective.
@@ II-B NIR-assited Image Restoration
Compared to single-image restoration, NIR images have the potential to assist in restoring details from degraded images. The earlier work [34] utilizes gradient constraints for NIR-assisted image denoising. Wang et al. [35] further improves the performance with deep learning methods. SSN [21] proposes a multi-task deep network with state synchronization modules. TC-GAN [36] fuses NIR images and RGB ones based on a texture conditional generative adversarial network. DCMAN [37] employs spatial-temporal-spectral priors to introduce NIR videos for low-light RGB video restoration. However, these methods have overlooked the color and structure inconsistency issues between the NIR images and RGB ones. CCDFuse [38] addresses this issue by combining the local modeling ability of convolutional blocks and the non-local modeling ability of transformer ones to extract local and global features of NIR and RGB images respectively. SANet [26] proposes a guided denoising framework by estimating a clean structure map for the noisy RGB image. Wan et al. [23] disentangle the color and structure components from the NIR images and RGB ones. Besides, a few works [39, 40, 22] incorporate different priors into the network design, like sparse coding [39], deep implicit prior [40] and deep inconsistency prior [22]. However, their intricate designs make it difficult to integrate them into existing advanced restoration networks, hindering their extensions and improvements.
@@ III Real-NAID: Real-World NIR-Assisted Image Denoising Dataset
Existing publicly available NAID datasets generally lack real-world noisy RGB images paired with clean RGB and NIR images, which limits the investigation in real-world NAID. Compared with the synthetic data, the noise in the real world is more complex and more difficult to remove, which makes it more challenging to fuse NIR and RGB images. Take synthetic DVD [22] dataset as an example, its noisy images are synthesized by simply adding Gaussian and Poisson noise to clean images, and the noise parameters are casually selected. Actually, this simple noise model is far from real-world complex noise. On the one hand, the noise parameters are Inaccurately calibrated. On the other hand, some image noise, e.g., row noise [44], quantization noise [44], and fixed pattern noise [45], are beyond the Gaussian and Poisson distribution. These noises are indispensable in low-light environments but are not considered.
To break such a limitation, we build a real-world NAID dataset, named Real-NAID. Specifically, we employ high ISO and short exposure time to capture the real-world noisy RGB images, as shown in Fig. 2 (a). For capturing the corresponding clean RGB images, we lower the ISO of the camera and appropriately increase the exposure time, as shown in Fig. 2 (b). To obtain paired NIR images, we activate NIR light to ensure a sufficient supply of NIR illumination and then capture them with a dedicated NIR camera, as shown in Fig. 2 (c). All images are captured with the Huawei X2381-VG surveillance camera, which is equipped with a built-in NIR illuminator specifically designed for capturing NIR images with spectral values between 750 nm and 1100 nm. To ensure image registration among multiple captures, we carefully position the camera and develop a remote control application to capture images of static objects. When capturing clean RGB images, the camera’s ISO is set to 600, as it does not introduce significant noise while allowing us to set a short exposure time to avoid blurry content. When capturing RGB images with low, middle, and high noise levels, we set ISO to 4000, 12000, and 32000, respectively. It is worth noting that we adjust the exposure time of each noise level to keep the brightness of noisy images relatively constant.
@@ IV-B Selective Fusion Module
Global Modulation Module. GMM should alleviate the adverse effects of color and luminance inconsistencies on NIR-RGB feature fusion. As shown in Fig. 4 (a), it takes the current NIR features $\mathbf{F_{N}}$ and the RGB ones $\mathbf{F_{R}}$ as inputs to estimate the NIR global modulation weights $\mathbf{W}^{g}_{\mathbf{N}}$ and the RGB ones $\mathbf{W}^{g}_{\mathbf{R}}$. Detailly, it first concatenates $\mathbf{F_{N}}\in\mathbb{R}^{C\times H\times W}$ and $\mathbf{F_{R}}\in\mathbb{R}^{C\times H\times W}$ as inputs, generating initial weights $\mathbf{W_{init}^{g}}\in\mathbb{R}^{2C\times H\times W}$ by a $1\times 1$ convolutional layer, two $1\times 1$ blocks and another $1\times 1$ convolutional layer. It can be written as,
@@
Local Modulation Module. The Local Modulation Module (LMM) should focus on the structure inconsistency between NIR images and RGB ones. We suggest increasing the receptive field to perceive more local structure information from a range of neighboring pixels, which is better for predicting the fusion weights of NIR and RGB features. In detail, LMM takes the globally modulated NIR features $\mathbf{F}^{g}_{\mathbf{N}}$ and the RGB ones $\mathbf{F}^{g}_{\mathbf{R}}$ as inputs to estimate the local NIR weights $\mathbf{W}^{l}_{\mathbf{N}}$ and the RGB ones $\mathbf{W}^{l}_{\mathbf{R}}$, as shown in Fig. 4 (a). Without complex network design, LMM is built upon GMM by replacing the $1\times 1$ convolutional layer in $1\times 1$ block to a large kernel depth-wise convolutional layer (DWConv) [48] for capturing more local information, as shown in Fig. 4 (c). Finally, $\mathbf{W}^{l}_{\mathbf{N}}$ and $\mathbf{W}^{l}_{\mathbf{R}}$ are employed to get the fused NIR and RGB feature $\mathbf{F_{NR}}$ as,
@@
Methods $\sigma=4$ $\sigma=8$FLOPs (G) Time (ms) #Params (M)
@@
Single-Image Denoising Uformer (CVPR’22) 29.58 / 0.8967 / 0.271 27.36 / 0.8632 / 0.352 19.16 1748 20.73
Restormer (CVPR’22) 29.67 / 0.9038 / 0.262 27.41 / 0.8741 / 0.343 70.59 2048 26.13
NAFNet (ECCV’22) 29.49 / 0.8959 / 0.263 27.29 / 0.8638 / 0.336 8.10 312 29.17
Xformer (ICLR’24) 29.57 / 0.8997 / 0.265 27.32 / 0.8690 / 0.347 73.31 5257 25.23
NIR-Assisted Image Denoising FGDNet (TMM’22) 23.91 / 0.8371 / 0.439 22.02 / 0.7374 / 0.436 38.67 479 1.63
SANet (CVPR’23) 27.68 / 0.8648 / 0.343 25.28 / 0.8304 / 0.413 161.06 2763 5.37
CUNet(TPAMI’20) 28.01 / 0.8558 / 0.332 26.07 / 0.8182 / 0.412 14.48 542 0.44
MNNet (IF’22) 28.48 / 0.8994 / 0.274 26.33 / 0.8697 / 0.353 23.68 1360 0.76
DVN (AAAI’22) 29.69 / 0.9062 / 0.236 27.43 / 0.8799 / 0.292 104.05 761 11.73
NIR-Uformer (Ours) 30.10 / 0.9188 / 0.192 28.03 / 0.9008 / 0.238 24.85 2500 24.51
NIR-Restormer (Ours) 30.22 / 0.9209 / 0.193 28.11 / 0.8701 / 0.260 89.17 2747 30.26
NIR-NAFNet (Ours) 30.08 / 0.9005 / 0.208 27.86 / 0.8664 / 0.273 13.17 462 35.06
NIR-Xformer (Ours) 30.29 / 0.9152 / 0.192 28.09 / 0.8947 / 0.230 95.63 7884 30.92
@@ V-A Experimental Settings
Implementation Details. We build our NIR-assisted denoising models by incorporating the proposed SFM into a CNN-based advanced denoising network (i.e., NAFNet [27]) and three Transformer-based ones (i.e., Uformer [11], Restormer [10] and Xformer [13]), which are dubbed NIR-NAFNet, NIR-Uformer, NIR-Restormer, and NIR-Xformer, respectively. Generally, based on the existing single image denoising network, we first add a new encoder branch for processing NIR images, and then utilize the proposed SFM to fuse the same scale NIR and RGB features. Finally, the fused features are fed into the corresponding decoder for reconstructing clean images. Network architectures are shown in Fig. 6, Fig. 7, Fig. 8 and Fig. 9. All models are trained by the Adam [49] optimizer with $\beta_{1}$ = 0.9 and $\beta_{2}$ = 0.999 for 120k iterations. The batch size is set to 32 and the patch size is set to $128\times 128$. For synthetic image denoising, the cosine annealing strategy [50] is employed to steadily decrease the learning rate from $2\times 10^{-4}$ to $1\times 10^{-6}$. For real-world image denoising, the initial learning rate is set to $3\times 10^{-4}$ and halved every 20k iterations. All experiments are conducted with PyTorch [51] on an Nvidia GeForce RTX A6000 GPU.
Figure 6: (a) Single image denoising with NAFNet [27]. (b) NIR-assisted image denoising with our NIR-NAFNet.
Figure 7: (a) Single image denoising with Restormer [10]. (b) NIR-assisted image denoising with our NIR-Restormer.
Figure 8: (a) Single image denoising with Uformer [11]. (b) NIR-assisted image denoising with our NIR-Uformer.
Figure 9: (a) Single image denoising with Xformer [13]. (b) NIR-assisted image denoising with our NIR-Xformer.
@@ V-B Comparison with State-of-the-Art Methods
Experiments are conducted by comparing our NIR-NAFNet, NIR-Uformer, NIR-Restormer, and NIR-Xformer with 9 models, including 4 single image denoising methods (i.e. NAFNet [27], Uformer [11], Restormer [10], and Xformer [13]) and 5 NIR-assisted image denoising methods (i.e. FGDNet [52], SANet [26], CUNet [39], MNNet [40], and DVN [22]). To evaluate the performance quantitatively, we calculate three metrics on the RGB channels, i.e., Peak Signal to Noise Ratio (PSNR), Structural Similarity (SSIM) [53] and Learned Perceptual Image Patch Similarity (LPIPS) [54]. The number of parameters (#Params), #FLOPs when processing a $128\times 128$ patch and the inference time when feeding a $1792\times 1008$ image are reported.
@@
Single-Image Denoising Uformer (CVPR’22) 25.56 / 0.7736 / 0.304 24.52 / 0.7418 / 0.347 23.31 / 0.7091 / 0.389
@@
NIR-Assisted Image Denoising FGDNet (TMM’22) 24.25 / 0.7676 / 0.368 22.89 / 0.7367 / 0.430 21.86 / 0.7080 / 0.509
@@ VI-A Comparisons with NAID Baselines
To demonstrate the effectiveness of our SFM, we compare our methods with NAID baselines using varying backbones. Specifically, we simply sum the NIR and RGB features as our baseline, as shown in Fig. 3 (b). We utilize three Transformer-based denoising networks (i.e., Uformer [11], Restormer [10], Xformer [13]) and a CNN-based one (i.e., NAFNet [27]) to conduct experiments, named ‘Uformer-Baseline’, ‘Restormer-Baseline’, ‘Xformer-Baseline’, and ‘NAFNet-Baseline’, respectively. The results are summarized in Table IV. First, the proposed SFM improves the performance of all NAID baselines with few computational cost additions. Second, the performance of the denoising model decreases as the noise level increases. This is because higher levels of noise cause more deterioration in RGB images, making image restoration more challenging. Third, the performance improvement from SFM becomes increasingly significant with higher noise levels. This is because the inconsistency between RGB and NIR images becomes more severe as the RGB image degradation becomes more severe. The simple summation in NAID baselines ignores it, whereas SFM better allocates fusion weights for RGB and NIR features.
@@
Methods Low-Level Noise Middle-Level Noise High-Level Noise #FLOPs (G) Time (ms) #Params (M)
@@
Spatial-Wised Transformer-Based Uformer-Baseline 25.80 / 0.7904 / 0.269 25.03 / 0.7687 / 0.294 24.02 / 0.7488 / 0.320 23.61 2284 23.39
NIR-Uformer 25.91 / 0.7917 / 0.276 25.14 / 0.7714 / 0.299 24.28 / 0.7534 / 0.321 24.85 2500 24.51
Channel-Wised Transformer-Based Restormer-Baseline 26.12 / 0.7948 / 0.265 25.38 / 0.7733 / 0.293 24.61 / 0.7587 / 0.316 86.75 2578 29.76
NIR-Restormer 26.22 / 0.7963 / 0.265 25.51 / 0.7767 / 0.293 24.76 / 0.7626 / 0.315 89.17 2747 30.26
Mixed Transformer-Based Xformer-Baseline 26.10 / 0.7927 / 0.264 25.23 / 0.7694 / 0.291 24.41 / 0.7553 / 0.314 90.78 7616 29.91
NIR-Xformer 26.19 / 0.7959 / 0.264 25.38 / 0.7729 / 0.293 24.66 / 0.7595 / 0.319 95.63 7884 30.92
CNN-Based NAFNet-Baseline 25.92 / 0.7883 / 0.277 25.00 / 0.7628 / 0.304 24.16 / 0.7429 / 0.329 11.72 462 34.16
NIR-NAFNet 26.06 / 0.7905 / 0.274 25.26 / 0.7676 / 0.303 24.48 / 0.7503 / 0.321 13.17 462 35.06